OpenStudy (anonymous):

A right circular conical vessel is being filled with green industrial waste at a rate of 75 cubic meters per second. How fast is the level rising after 250π cubic meters have been poured in? The cone has a height of 70 m and a radius 30 m at its brim. (The volume of a cone of height h and cross-sectional radius r at its brim is given by V = 1 3 πr2h. Round your answer to two decimal places.)

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (dumbcow):

the cone of water will be proportional to conical vessel \[\frac{r}{30} = \frac{h}{70}\] \[r = \frac{3}{7} h\] now define volume in terms of height(water level) \[V = \frac{1}{3}\pi (\frac{3}{7}h)^{2}h = \frac{3}{49}\pi h^{3}\] now using chain rule \[\frac{dV}{dt} = \frac{dV}{dh}*\frac{dh}{dt}\] given dV/dt = 75, solve for dh/dt take derivative to find dV/dh also find water level "h" when V = 250pi